OpenStudy (anonymous):

Some high school students conducted an experiment as part of their school project to understand the relationship between the height a parasail can attain and the speed of the motorboat that tows it. The data table lists their observation about the height of the parasail at different speeds. Which equation represents the height of the parasail (y) based on the speed of the motorboat (x)? Speed of the Motorboat (in miles per hour) Height of the Parasail from the Water Surface (in feet) 12.5 5 14.5 10 16.5 15 18.5 20 20.5 25 22.5 30

OpenStudy (anonymous):

These are the option: A y=2.5x B y=0.4x C y=2.5x-26.25 D y=0.4+4.2 E y=0.4-4.2